ESPERANTOLOGIO

ESPERANTO STUDIES

EES

Kajero 1  1999  Issue No. 1

Redaktas Editor

Christer Kiselman

Redakta komitato Editorial Board

Detlev Blanke José Borges Neto Probal Dasgupta Michel Duc Goninaz Gotoo Hitosi Edmund Grimley Evans Koutny Ilona Erich-Dieter Krause Sergej Kuznecov Jouko Lindstedt Liu Haitao Ingrid Maier Geraldo Mattos Marc van Oostendorp Sergio Pokrovskij Humphrey Tonkin John C. Wells Yamasaki Seikô

Varna

Upsala Universitato Uppsala University